Jack has a 75 foot vertical TV antenna on his roof. He wants to connect a guy wire from the top of the antenna to a hook on the roof. The hook is located 15 feet from the base of the antenna. Ho many feet of wire will be needed?
Answer by checkley77(12844)

You have a right triangle.
75^2+15^2=c^2
5,625+225=c^2
5,850=c^2
c^2=5,850
c=sqrt5,850
c=76.4853 ft. ans.